# Reinforcement Learning for Ballbot Navigation in Uneven Terrain

### v1

@@ Abstract
Ballbot (i.e. Ball balancing robot) navigation usually relies on methods rooted in control theory (CT), and works that apply Reinforcement learning (RL) to the problem remain rare while generally being limited to specific subtasks (e.g. balance recovery). Unlike CT based methods, RL does not require (simplifying) assumptions about environment dynamics (e.g. the absence of slippage between the ball and the floor). In addition to this increased accuracy in modeling, RL agents can easily be conditioned on additional observations such as depth-maps without the need for explicit formulations from first principles, leading to increased adaptivity. Despite those advantages, there has been little to no investigation into the capabilities, data-efficiency and limitations of RL based methods for ballbot control and navigation. Furthermore, there is a notable absence of an open-source, RL-friendly simulator for this task. In this paper, we present an open-source ballbot simulation based on MuJoCo, and show that with appropriate conditioning on exteroceptive observations as well as reward shaping, policies learned by classical model-free RL methods are capable of effectively navigating through randomly generated uneven terrain, using a reasonable amount of data (four to five hours on a system operating at $500$hz). Our code is available at https://github.com/salehiac/OpenBallBot-RL.
@@ I Introduction
The term Ballbot (i.e. Ball balancing robot) refers to a family of dynamically stable, underactuated robots that are similar in principle to an inverted pendulum mounted on a spherical base [1, 2, 3]. The single spherical base results in omnidirectional mobility and agility/maneuverability, especially in comparison with more conventional, statically stable wheeled robots. However, ballbot control is notoriously challenging due to several factors such as the system’s high non-linearity and sensitivity to perturbations [3, 4, 5, 6]. Almost all approaches to Ballbot control in the literature are based on methods from control theory that require a mathematical model of the ballbot. Those models generally make several simplifying assumptions [7, 8, 6], , e.g. that the ball moves in a perfectly horizontal plane, that there is no slippage between the floor and the ball and that there is no deformation.
@@ II-A The ballbot
Ballbot control methods in general define $q\triangleq[x,y,\theta_{0},\theta_{1},\theta_{2}]^{T}$ (where $[x,y]^{T}$ is the planar position and the $\theta_{i}$ specify the body’s orientation) and use the Euler-Lagrange Equation
@@ II-B Problem formulation
Let $<\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{P}_{0},\mathcal{R},\gamma>$ denote a Markov Decision Process (MDP) with $\mathcal{S},\mathcal{A}$ respectively the set of states and actions, $\mathcal{P}(s^{\prime}|s,a)$ the transition probabilities, $\mathcal{P}_{0}$ the initial state distribution, $\mathcal{R}(s,a,s^{\prime})$ the reward function, and $\gamma\in[0,1)$ the discount factor. Let us also denote $J(\theta)=E_{\tau\sim p_{\theta}}[\sum_{t}^{\infty}r_{t}\gamma^{t}]$ the discounted cumulative rewards with the expectation taken over trajectories $\tau$ sampled from the policy $\pi_{\theta}\triangleq p_{\theta}(a|s)$.
@@ III Proposed solution
A ballbot’s control an proprioception usually operate at a high frequency $f_{high}$ that can not be matched by most camera systems. To provide unambiguous observations at the same frequency as $f_{high}$, we use the most recent depth images $z^{1}_{t^{\prime}},z^{2}_{t^{\prime}}$ at each timestep and append the time difference $\Delta_{t}=(t-t^{\prime})$ to the observations. The full observation vector is then defined by
@@
$\mathcal{R}(s_{t},a_{t})=\alpha_{1}v_{t}^{T}g+\mathbf{1}_{\{s_{t}\notin S_{F}\}}\alpha_{2}-\alpha_{3}||a_{t}||^{2}$ (4)
@@
Figure 4: The policy used to navigate uneven terrain. Two low resolution $128\times 128$ depth images, one from each depth camera, are fed to a pretrained encoder that maps each of them to an embedding in $\mathbb{R}^{20}$. These embeddings are then concatenated to the proprioceptive observations: orientation, angular velocity, body velocity, angular velocities of each omniwheel (the latter are in local coordinates), as well as the last command vector sent at the previous timestep. To avoid state ambiguities arising from the differences in observation frequencies ($500$hz for proprioceptive readings vs $\sim 80hz$ for the depth cameras), we also concatenate the time $\Delta t$ that has elapsed since the last depth image observation was received. This $56$ dimensional vector is then fed into a small MLP which is trained to predict torque commands that are sent to the omniwheel motors.
@@ IV Simulated environment
Figure 5 (left) shows our simulated ballbot, including the two $128\times 128$ depth cameras that are oriented towards the contact point between the ball and the ground. As shown in figure 5 (right), omniwheels are modeled as capsules with anisotropic friction. Note that as of this writing, The current stable MuJoCo release requires a small modification to allow anisotropic friction to function as intended. This tweak is provided as a patch in our public repository, and was suggested by the developers of MuJoCo [^2] See the discussion at
https://github.com/google-deepmind/mujoco/discussions/2517
Note that robot control is performed at $500hz$, and while proprioceptive observations $o^{pro}_{t}$ are read at that same frequency, the depth cameras operate at $\sim 80hz$. As discussed in section §III, unambiguous observations $o_{t}$ are provided at $500hz$ by concatenating the relative timestamp $\Delta t$ to the depth image encodings and proprioceptive readings composing the state.
The distribution $\mathcal{P}_{perlin}$ is based on the snoise2 function provided by [20].
@@ V Experiments
In this section, we show that our trained policies have effectively learned to navigate in uneven terrain and generalize to previously unseen situations. We also provide some qualitative results, more of which can be found in the project’s public repository. We note that since ballbot navigation in uneven terrain is an extremely under-explored area, we are not aware of any existing method that would serve as a reasonable comparative baseline.
During training, the maximum length of each episode was fixed to a horizon of $H_{train}\triangleq 4000$ steps, with each episode ending either in case of failure (i.e. the robot falling down) or after reaching $H_{train}$ steps. We conducted five training experiments, each with a total training budget of $68e$ 6 environment steps (one step taking $\sim 2ms$ in simulation time) and a different RNG seed.
The generalization performance of the policy was evaluated at regular intervals, using $N=10$ randomly sampled, previously unseen terrains. The results are reported in figure 6. In each training experiment, the policy has learned to robustly balance the robot in about $63e$ 6 timesteps. This is evident from the fact that the average length of evaluation episodes (figure 6, bottom) starts to plateau at $\sim 4000$ environment steps, which is the maximum that is reachable in our settings (shorter episodes would mean the robot has fallen). At this time, the policy has primarily learned to optimize for the second and third terms of equation 4. During the remainder of each experiment (environment timesteps $\sim 3e6$ to $68e$ 6), the remaining reward term (goal direction) is optimized, until the reward plateaus slightly below a value of $86.0$. Noting $r_{max}^{eval}(i)$ the maximum of average evaluation rewards reached during an experiment $i$, $\texttt{median}(\{r_{max}^{eval}(i)\}_{i=1}^{5})\approx 85.5$, which given our choice of hyperparameters for the reward function (appendix B), translates to an average speed of $\sim 0.5m/s$ in the desired goal direction.
An example of policy behavior in a previously unseen terrain is given in figure 8 for qualitative illustration. More qualitative examples are available in our repository.
@@ VI Discussion
Up to this point, we have demonstrated that with appropriate reward shaping and a sufficiently informative observation space, model-free RL algorithms, in particular PPO, can learn policies that allow the ballbot to navigate in uneven terrain and that generalize well to unseen situations. Interestingly, the $68e$ 6 transitions that were used to obtain those results are equivalent to $\sim 4.4$ hours of data, which would be reasonable on a real system. Notice however that we have not aimed for data-efficiency, and that this aspect is likely to be improved by using other actor-critic methods (e.g. SAC) or model-based methods such as DreamerV3[19].
@@ VII Related work
Several ballbot simulation environments have been developed by different authors (e.g. [21, 7, 22, 23], however, to our knowledge no functional open-source ballbot simulations have been published [^3] A few repositories exists on github, but seem to have been abandoned mid-development.. Furthermore, simulations developed as part of prior research projects often rely on physics engines that, unlike MuJoCo, are not idea for RL research, mainly due to their computational costs and the difficulty of running parallel environments.
@@ Appendix A Network architecture
The encoder is a vanilla CNN whose details are given in table I. The $56$ dimensional feature vector resulting from concatenating proprioceptive observations and depth embeddings (see figure 4) are then passed into a small MLP with $5$ linear layers with hidden dimension $128$, LeakyReLu and no normalization.
@@
Conv1 input_c=1, output_c=32, kernel size=3, stride=2, padding=1
@@
Conv2 input_c=32, output_c=32, kernel size=3, stride=2, padding=1
@@ Appendix B Training details
In our implementation, the coefficient from the reward function (equation 4) where set to $\alpha_{1}=0.01,\alpha_{2}=0.02,\alpha_{3}=0.0001$.
We used the PPO implementation from stable_baselines3 [27]. The hyperparameters used for training are given in table III. As is usual with PPO, we found that tuning the clip range, entropy coefficient, epochs per updates as well as LR scheduling were critical for training success. Starting with and LR of $41e-$ 4, we manually scheduled the LR by dividing it by constant factors after a specific number of environment timesteps were reached.

### v2

@@ Abstract
Ballbot (Ball balancing robot) navigation relies on methods rooted in control theory (CT), and works that apply Reinforcement Learning (RL) to the problem remain rare while being limited to specific subtasks. Unlike CT based methods, RL does not require simplifying assumptions about environment dynamics. In addition to this increased accuracy in modeling, RL agents can easily be conditioned on observations such as depth-maps without the need for explicit formulations from first principles, leading to increased adaptivity. Despite those advantages, there has been little to no investigation into the capabilities, data-efficiency and limitations of RL for ballbot control and navigation. Furthermore, there is a notable absence of an open-source, RL-friendly simulator for this task. We present an open-source ballbot simulation based on MuJoCo, and show that with appropriate conditioning and reward shaping, classical model-free RL methods can produce policies that effectively navigate randomly generated uneven terrain, using a reasonable amount of data (four to five hours on a system operating at 500hz). To our knowledge, none of the current CT based methods are able to tackle such scenarios. Our code is available at https://github.com/salehiac/OpenBallBot-RL.
@@ I Introduction
The term Ballbot (i.e. Ball balancing robot) refers to a family of dynamically stable, underactuated robots that are similar in principle to an inverted pendulum mounted on a spherical base [15, 13, 4]. The single spherical base results in omnidirectional mobility and agility/maneuverability, especially in comparison with more conventional, statically stable wheeled robots. However, ballbot control is notoriously challenging due to several factors such as the system’s high non-linearity and sensitivity to perturbations [4, 5, 25, 18]. Almost all approaches to Ballbot control in the literature are based on methods from control theory that require a mathematical model of the ballbot. Those models generally make several simplifying assumptions [27, 10, 18], e.g. that the ball moves in a perfectly horizontal plane, that there is no slippage between the floor and the ball and that there is no deformation.
@@ II-A The ballbot
Ballbot control methods in general define $q\triangleq[x,y,\theta_{0},\theta_{1},\theta_{2}]^{T}$ (where $x,y]^{T}$ is the planar position and the $\theta_{i}$ specify the body’s orientation) and use the Euler-Lagrange Equation
@@ II-B Problem formulation
Let $\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{P}_{0},\mathcal{R},\gamma>$ denote a Markov Decision Process (MDP) with $\mathcal{S},\mathcal{A}$ respectively the set of states and actions, $\mathcal{P}(s^{\prime}|s,a)$ the transition probabilities, $\mathcal{P}_{0}$ the initial state distribution, $\mathcal{R}(s,a,s^{\prime})$ the reward function, and $\gamma\in[0,1)$ the discount factor. Let us also denote $J(\theta)=E_{\tau\sim p_{\theta}}[\sum_{t}^{\infty}r_{t}\gamma^{t}]$ the discounted cumulative rewards with the expectation taken over trajectories $\tau$ sampled from the policy $\pi_{\theta}\triangleq p_{\theta}(a|s)$.
@@ III Proposed solution
A ballbot’s control and proprioception usually operate at a high frequency $f_{high}$ that can not be matched by most camera systems. To provide unambiguous observations at the same frequency as $f_{high}$, we use the most recent depth images $z^{1}_{t^{\prime}},z^{2}_{t^{\prime}}$ at each timestep and append the time difference $\Delta_{t}=(t-t^{\prime})$ to the observations. The full observation vector is then defined by
@@
[…]
@@
Figure 4: The policy used to navigate uneven terrain. Two low resolution $128\times 128$ depth images, one from each depth camera, are fed to a pretrained encoder that maps each of them to an embedding in $\mathbb{R}^{20}$. These embeddings are then concatenated to the proprioceptive observations: orientation, angular velocity, body velocity, angular velocities of each omniwheel (the latter are in local coordinates), as well as the last command vector sent at the previous timestep. To avoid state ambiguities arising from the differences in observation frequencies (500500hz for proprioceptive readings vs $\sim 80hz$ for the depth cameras), we also concatenate the time $\Delta t$ that has elapsed since the last depth image observation was received. This $56$ dimensional vector is then fed into a small MLP which is trained to predict torque commands that are sent to the omniwheel motors.
@@ IV Simulated environment
Figure 5 (left) shows our simulated ballbot, including the two $128\times 128$ depth cameras that are oriented towards the contact point between the ball and the ground. As shown in figure 5 (right), omniwheels are modeled as capsules with anisotropic friction. Note that as of this writing, The current stable MuJoCo release requires a small modification to allow anisotropic friction to function as intended. This tweak is provided as a patch in our public repository, and was suggested by the developers of MuJoCo [^2] See the discussion at https://github.com/google-deepmind/mujoco/discussions/2517.
Note that robot control is performed at $500hz$, and while proprioceptive observations $o^{prop}_{t}$ are read at that same frequency, the depth cameras operate at $\sim 80hz$. As discussed in section §III, unambiguous observations $o_{t}$ are provided at $500hz$ by concatenating the relative timestamp $\Delta t$ to the depth image encodings and proprioceptive readings composing the state.
The distribution $\mathcal{P}_{perlin}$ is based on the snoise2 function provided by [3] [^3] We set its parameters to scale=25.0,octaves=4,persistence=0.2,
lacunarity=2.0 in this work.
@@ V Experiments
In this section, we:
  • show that our trained policies have effectively learned to navigate in uneven terrain and generalize to unseen situations;
  • provide some qualitative results (more of which can be found in the project’s public repository); and
  • use comparisons with a classical baseline from control theory to assess the difficulty of the uneven terrains from $\mathcal{P}_{0}$.
Regarding the last point, we emphasize that ballbot navigation in uneven terrain is an extremely under-explored area, and that we are not aware of any existing method that would serve as a reasonable comparative baseline. As discussed in previous sections (e.g. §II), control theory based methods, aside from limiting observations to proprioceptive ones, make the assumption that motion is perfectly planar. As a result, any comparison between those approaches and ours will be unfair. That being said, the performance of algorithms based on control theory can be used to assess the difficulty of the terrain distribution $\mathcal{P}_{0}$ and verify that they are not too similar to the flat terrain case.
Learning and generalization. During training, the maximum length of each episode was fixed to a horizon of $H_{train}\triangleq 4000$ steps, with each episode ending either in case of failure (i.e. the robot falling down) or after reaching $H_{train}$ steps. We conducted five training experiments, each with a total training budget of $8e6$ environment steps (one step taking $\sim 2ms$ in simulation time) and a different RNG seed.
@@
Figure 9: Comparison between cascaded PIDs and a trained policy, on a horizon of $4000$ timesteps.
The generalization performance of the policy was evaluated at regular intervals, using $N=10$ randomly sampled, previously unseen terrains. The results are reported in figure 6. In each training experiment, the policy has learned to robustly balance the robot in about $3e6$ timesteps. This is evident from the fact that the average length of evaluation episodes (figure 6, bottom) starts to plateau at $\sim 4000$ environment steps, which is the maximum that is reachable in our settings (shorter episodes would mean the robot has fallen). At this time, the policy has primarily learned to optimize for the second and third terms of equation 4. During the remainder of each experiment (environment timesteps $\sim 3e6$ to $8e6$), the remaining reward term (goal direction) is optimized, until the reward plateaus slightly below a value of $86.0$. Noting $r_{max}^{eval}(i)$ the maximum of average evaluation rewards reached during an experiment $i$, $\texttt{median}(\{r_{max}^{eval}(i)\}_{i=1}^{5})\approx 85.5$, which given our choice of hyperparameters for the reward function (appendix B), translates to an average speed of $\sim 0.5m/s$ in the desired goal direction.
Qualitative results. An example of policy behavior in a previously unseen terrain is given in figure 8 for qualitative illustration. More qualitative examples are available in our repository.
Terrain difficulty. We assessed the difficulty of the terrains sampled from $\mathcal{P}_{0}$ by observing the performance of cascaded PIDs, which can be considered as a standard baseline for ballbot control [14, 4]. As is usual, the inner loop of our cascaded PIDs was responsible for balancing, while the outer loop (producing the roll and pitch setpoints for the inner loop) was responsible for reaching the target velocity in the $0,1]^{T}$ direction. The cascaded PIDs were tuned for flat terrain and we compared them to a policy trained on uneven terrain in both flat and uneven environments. The results, which are based on a number of $100$ randomly sampled terrains and a horizon of $4000$, are summarized in figure 9. Note that only velocity rewards (i.e. first term of equation 4) are taken into account for these comparisons. As shown in figure 9 (right), in flat terrain, both policies balance the robot for the maximum possible horizon, while the cascaded PIDs reach higher velocity rewards than our trained policy (figure 9 (left)). We hypothesize that this is because the trained policy has not been trained on flat terrain, and that it therefore moves in a more conservative way due to the risk of losing balance in uneven terrain.
However, the performance of cascaded PIDs plummets on uneven terrain: while the robot is able to keep its balance for $\sim 3000$ timesteps on average —with considerable standard deviation —as shown in figure 9 (right), it is unable to navigate in the desired direction (figure 9 (left)), often navigating in random directions, reaching an almost null reward on average, while only collecting negative rewards in some episodes (as illustrated by the standard deviation in figure 9 (left)). This indicates that the terrains sampled from $\mathcal{P}_{0}$ are not too close to the planar case, and that navigating them is indeed challenging. As shown in both plots from figure 9, our trained policy is only slightly less effective in uneven terrains from $\mathcal{P}_{0}$ than in flat terrain.
@@ VI Discussion
Up to this point, we have demonstrated that with appropriate reward shaping and a sufficiently informative observation space, model-free RL algorithms, in particular PPO, can learn policies that allow the ballbot to navigate in uneven terrain and that generalize well to unseen situations. Interestingly, the $8e6$ transitions that were used to obtain those results are equivalent to $\sim 4.4$ hours of data, which would be reasonable on a real system. Notice however that we have not aimed for data-efficiency, and that this aspect is likely to be improved by using other actor-critic methods (e.g. SAC) or model-based methods such as DreamerV3[8].
@@ VII Related work
Several ballbot simulation environments have been developed by different authors (e.g. [9, 27, 23, 16], however, to our knowledge no functional open-source ballbot simulations have been published [^4] A few repositories exists on github, but seem to have been abandoned mid-development.. Furthermore, simulations developed as part of prior research projects often rely on physics engines that, unlike MuJoCo, are not ideal for RL research, mainly due to their computational costs and the difficulty of running parallel environments.
@@ Appendix A Network architecture
The encoder is a vanilla CNN whose details are given in table I. The $56$ dimensional feature vector resulting from concatenating proprioceptive observations and depth embeddings (figure 4) is then passed into a small MLP with $5$ linear layers with hidden dimension $128$, LeakyReLu and no normalization.
@@
Conv1 in=1, out=32, kernel sz=3, stride=2, padding=1
@@
Conv2 in=32, out=32, kernel sz=3, stride=2, padding=1
@@ Appendix B Training details
The coefficient from the reward function (equation 4) were set to $\alpha_{1}=0.01,\alpha_{2}=0.02,\alpha_{3}=0.0001$. We used the PPO implementation from stable_baselines3 [19]. The hyperparameters used for training are given in table II. As is usual with PPO, we found that tuning the clip range, entropy coefficient and epochs per updates was critical. Starting with a LR of 1e41e-4, we manually scheduled the LR by dividing it by constant factors after a specific number of environment timesteps were reached.